The following condensed information was reported by Peabody Toys, Inc., for 2018 and 2017: (S in 000s) 2018 2017 Income statement information Net sales Net income Balance sheet information Current assets Property, plant, and equipment (net) $5,800 $4,800 136 184 $ 860 810 1,640 1,290 $2,500 $2,100 Total assets Current liabilities Long-term liabilities Paid-in capital Retained earnings 830 810 300 160 $1,180 810 300 210 Liabilities and shareholders equity $2,500 $2,100 Required: 1. Determine the following ratios for 2018: (Round your percentage answers to 1 decimal place.) 2. Determine the amount of dividends paid to shareholders during 2018. (Enter your answer in dollars. 1a. Profit margin on sales 1b. Return on assets 1c. Return on shareholders equity 2. Dividends paid
Answer #1

1a.Profit Margin on Sales = Net income/Net Sales

= 184/5,800

= 3.2%

1b.Return on Assets = Net income/Average Assets

Average Assets = (2,500+2,100)/2 = 2,300

Return on Assets= 184/2,300

= 8

1c.Return on Shareholder’s Equity = Net Income/Shareholders Funds

Shareholders Equity = paid in capital + retained earnings

Average = (510+460)/2

= 485

Return on Equity = 184/485

= 37.9%

2.Net Income for the year = $184

Increase in Retained Earnings = $50

Dividend paid = $134